### v2

@@ Acknowledgements
This research was supported by the National University of Singapore and benefited from access to computational resources provided by CogAI4S. We are deeply grateful to the human annotators who contributed their time and expertise to this project, including Yiming Tang, Qiran Zou, Arash Lagzian, Samson Yu, Ye Zhang and Yingtao Zhu. Their careful assessments played a vital role in the validation of the LanSE method. Thanks to Xuming Ran for his valuable advice and guidance during the early beginning of the project. We are also grateful to Vaishnavh Nagarajan for his insightful feedback and constructive suggestions that helped shape the direction of this work.
A. Qualitative Examples of LanSE Neurons
In this section, we provide qualitative examples of the interpretable visual patterns used in our evaluation method, LanSE. These visualizations contain individual visual patterns that encodes meaningful explanations and the set of images that can activate them, including both photos and generated images.
Figure 14: Visual pattern examples for the human subgroup of semantic meaning.
Figure 15: Visual pattern examples for the animal subgroup of semantic meaning.
Figure 16: Visual pattern examples for the object subgroup of semantic meaning.
Figure 17: Visual pattern examples for the activity subgroup of semantic meaning.
Figure 18: Visual pattern examples for the environment subgroup of semantic meaning.
Figure 19: Visual pattern examples for the style subgroup of visual realism.
Figure 20: Visual pattern examples for the artifact subgroup of visual realism.
Figure 21: Visual pattern examples for the distortion subgroup of physical plausibility.
Figure 22: Visual pattern examples for the structure subgroup of physical plausibility.
Figure 23: Visual pattern examples for the CXR-LanSE built on chest X-ray images.
B. Qualitative Examples of Image-caption Pairs and Activated Neurons
Figure 24: Image example and the LanSE neurons it highly activates.
Figure 25: Image example and the LanSE neurons it highly activates.
Figure 26: Image example and the LanSE neurons it highly activates.
Figure 27: Image example and the LanSE neurons it highly activates.
C. LMM Analysis Prompts
@@
D. Method Extention
Certain failure modes in generative models, such as physics violations (e.g., distorted limbs, extra fingers), are both rare and critical, yet difficult to discover through random sampling. To address this, we develop an interactive human-in-the-loop framework that enables targeted visual pattern discovery at scale while reducing human annotation cost, as shown in Figure 28.
@@ D. Method Extention
Figure 28: The general pipeline of our extended method used for discovering monosemantic neurons that selectively respond to targeted semantics (physics violations).
E. Mathematical Definitions of the Four Metrics
@@
F. Human Evaluation Protocols
In this section, we detail the procedures we conduct for the human evaluations. Our human evaluation protocol for LansE’s visual pattern detection accuracy is shown in Figure 29. Human annotators are trained and tasked to identify if the image generally matches with the interpretation from LanSE. We release the human evaluation website for transparency: https://512d-223-109-239-18.ngrok-free.app/.
Figure 29: A demonstration of our human evaluation web interface. Human annotators are tasked to identify if the image generally matches with the LanSE interpretations.
G. The Choice of LanSE Neuron Groups
@@ H. Ablation Studies
Metrics v.s. the Activation Threshold $\tau$ We tune the activation threshold used for the LanSE neurons and collect all the diagnostic metrics. The results are listed in Figure 31 and Figure 30.
Figure 30: LanSE metrics varying according to the threshold choices for two models. Higher thresholds mean the neurons are less likely activated and leads to smaller metric values.
Figure 31: LanSE metrics varying according to the threshold choices for two groups. Higher thresholds mean the neurons are less likely activated and leads to smaller metric values. Lower metrics gives better contrast between models, making it easier to distinguish them.
Accuracies v.s. the Activation Threshold $\tau$ We tune the activation threshold used for the LanSE neurons and utilize an LMM to evaluate the ratio of images that can activate the neurons generally match the natural language interpretations. The results are listed in Figure 32.
Figure 32: LanSE’s signal accuracies varying according to the threshold choices. Higher thresholds means stricter constraints on the identification of visual pattern and therefore higher accuracies for the signals.

### v3

@@ Acknowledgements
We are deeply grateful for the computational resources provided by the Cognitive AI for Science lab (https://www.cogai4sci.com/). Great thanks to the human annotators who contributed their time and expertise to this project, including Yiming Tang, Qiran Zou, Arash Lagzian, Samson Yu, Ye Zhang and Yingtao Zhu. Thanks to Xuming Ran for his valuable advice and guidance during the early beginning of the project. We are also grateful to Vaishnavh Nagarajan for his insightful feedback and constructive suggestions that helped shape the direction of this work. Finally, we acknowledge the use of Claude (Anthropic) for language editing and improving readability.
Funding
This work was supported by Singapore MOE Tier 2 grant T2EP20125-0026 and Singapore MOH SIMFONI grant to D.L.
Declarations
Author Contributions
Yiming T. and D.L. conceived the project. Yiming T. developed the methodology, conducted experiments, and performed data analysis. A.L. and Q.Z. contributed to experimental validation and manuscript drafting. Yingtao Z. and Ye Z. assisted with experimental validation. S.A., T.N., Yih-Chung T., E.A., C.C., and Y.D. provided critical feedback and helped conceive key experiments. D.L. supervised the project. All authors contributed to manuscript preparation and approved the final version.
Competing Interests
The authors declare no competing interests.
Appendix A: Qualitative Examples of LanSE Neurons
In this section, we provide qualitative examples of the interpretable visual patterns used in our evaluation method, LanSE (Figs. A1A10). These visualizations contain individual visual patterns that encodes meaningful explanations and the set of images that can activate them, including both photos and generated images.
Figure A1: Visual pattern examples for the human subgroup of semantic meaning.
Figure A2: Visual pattern examples for the animal subgroup of semantic meaning.
Figure A3: Visual pattern examples for the object subgroup of semantic meaning.
Figure A4: Visual pattern examples for the activity subgroup of semantic meaning.
Figure A5: Visual pattern examples for the environment subgroup of semantic meaning.
Figure A6: Visual pattern examples for the style subgroup of visual realism.
Figure A7: Visual pattern examples for the artifact subgroup of visual realism.
Figure A8: Visual pattern examples for the distortion subgroup of physical plausibility.
Figure A9: Visual pattern examples for the structure subgroup of physical plausibility.
Figure A10: Visual pattern examples for the CXR-LanSE built on chest X-ray images.
Appendix B. Qualitative Examples of Image-caption Pairs and Activated Neurons
Figure A11: Image example and the LanSE neurons it highly activates.
Figure A12: Image example and the LanSE neurons it highly activates.
Figure A13: Image example and the LanSE neurons it highly activates.
Figure A14: Image example and the LanSE neurons it highly activates.
Appendix C. LMM Analysis Prompts
@@
Appendix D. Method Extention
Certain failure modes in generative models, such as physics violations (e.g., distorted limbs, extra fingers), are both rare and critical, yet difficult to discover through random sampling. To address this, we develop an interactive human-in-the-loop framework that enables targeted visual pattern discovery at scale while reducing human annotation cost, as shown in Figure A15.
@@ Appendix D. Method Extention
Figure A15: The general pipeline of our extended method used for discovering monosemantic neurons that selectively respond to targeted semantics (physics violations).
Appendix E. Mathematical Definitions of the Four Metrics
@@
Appendix F. Human Evaluation Protocols
In this section, we detail the procedures we conduct for the human evaluations. Our human evaluation protocol for LansE’s visual pattern detection accuracy is shown in Figure A16. Human annotators are trained and tasked to identify if the image generally matches with the interpretation from LanSE. We release the human evaluation website for transparency: https://512d-223-109-239-18.ngrok-free.app/.
Figure A16: A demonstration of our human evaluation web interface. Human annotators are tasked to identify if the image generally matches with the LanSE interpretations.
Appendix G. The Choice of LanSE Neuron Groups
@@ Appendix H. Ablation Studies
Metrics v.s. the Activation Threshold $\tau$ We tune the activation threshold used for the LanSE neurons and collect all the diagnostic metrics. The results are listed in Figure A18 and Figure A17.
Figure A17: LanSE metrics varying according to the threshold choices for two models. Higher thresholds mean the neurons are less likely activated and leads to smaller metric values.
Figure A18: LanSE metrics varying according to the threshold choices for two groups. Higher thresholds mean the neurons are less likely activated and leads to smaller metric values. Lower metrics gives better contrast between models, making it easier to distinguish them.
Accuracies v.s. the Activation Threshold $\tau$ We tune the activation threshold used for the LanSE neurons and utilize an LMM to evaluate the ratio of images that can activate the neurons generally match the natural language interpretations. The results are listed in Figure A19.
Figure A19: LanSE’s signal accuracies varying according to the threshold choices. Higher thresholds means stricter constraints on the identification of visual pattern and therefore higher accuracies for the signals.